OpenStudy (anonymous):

(b-4)/6 = b/2 solve for b?

OpenStudy (alexandervonhumboldt2):

first multiply both sides by 6. (b-4)/6*6=b/2*6 b-4=b*3 b-4=3b subtract b from both sides: b-b-4=3b-b -4=2b divide both sides by 2: -4/2=2b/2 b=-2